OpenStudy (anonymous):

Write an algebra1. The length of a rectangle is 3 more than twice its width. The perimeter is 48 feet. 2. Fifteen more than four times a number is 6 more than five times the number. Use "n" for the number. 3. Martha and Sally had lunch together on Tuesday. Sally paid $2 more than Martha. The total bill was $5.64. 4. Bob paid $3 less than Tim at a local pizza parlor. Together they spent $7.80. 5. There were six more women on a committee than men. The committee was made up of 30 people. 6. Three consecutive integers have the sum of 153.

OpenStudy (anonymous):

i just need 1 and 6

OpenStudy (anonymous):

1. Length is 3 more than twice the width. Call lethth 'L' and width "w" then 'twice the width' is 2w and three more than that is +3 so: L=2w+3

OpenStudy (anonymous):

Permeter is 48 and P = 2L + 2w. so 2L+2W=48. From there you can substitute one into the other and solve.

OpenStudy (anonymous):

6. 'Three conssecutive numbers' --> n, n+1, n+2 so n+(n+1)+(n+2)=153 3n+3=153 3n=150 n= 150/3 n=50

OpenStudy (anonymous):

Five times the first of three consecutive even integers is fourteen more than three times the second. What is the third integer? @twvogels

OpenStudy (anonymous):

(5*n) +14 = 3*(n+2) 5n+14=3n+6 8=-2n n=-4 Therefore the numbers are -4, -2, and 0, the third one is 0.

OpenStudy (anonymous):

To check: (5*-4)+14 = -6 and 3*(-4+2) =-6 so it checks.

OpenStudy (anonymous):

thanks